OpenStudy (anonymous):

A furniture company charges a base rate plus a rate per day for an office furniture set. There are two available plans. The Super Saver Plan charges $100 +$5 per day. The Best Deal Plan charges $150 +$3 per day. How many days make the Super Plan more expensive than the Best Deal Plan?

OpenStudy (radar):

I would find out how many days it would take for them become equal, then the next day it would be more expensive

OpenStudy (radar):

Let x = number of days that both plans would cost the same (equal) First do The Super Plan: Cost would be, $100+$5x Now do the Best Deal Plan: Cost would be, $150 +$3x

OpenStudy (radar):

I am going to drop the $ sign 100x+5x=150+3x Setting the two plans to equal each other then solve for X

OpenStudy (radar):

redo 100+5x=150+3x NO 100x lol 2x=50 then x=25 25 days they are equal, 26 days The Best Deal becomes the best deal.